OpenStudy (anonymous):

A playground is surrounded by a fence 2.0 m high. A boy who plays in the playground kicks a ball vertically upward from a height of 0.2 m with a speed of 10 m/s. How much time does the ball spend above the height of the fence?

OpenStudy (anonymous):

find the solutions of \[-16t^2+10t+.2=2\] or \[-16t^2+10t-1.8=0\] this will give you the time it first reaches the top of the fence and the time it reaches it again. then subtract the smaller one from the bigger one to get the time between them

OpenStudy (anonymous):

might be easier to solve \[-160t^2+100t-18=0\] and if fact since you only want the difference of the two roots you can use \[\frac{\sqrt{b^2-4ac}}{2}\]

OpenStudy (anonymous):

typo there should be \[\frac{\sqrt{b^2-4ac}}{a}\]

OpenStudy (anonymous):

-16t stand for what ?

OpenStudy (anonymous):

damn! meters not feet!!!

OpenStudy (anonymous):

ok scratch that. solve \[-4.9t^2+10t-1.8=0\]

OpenStudy (anonymous):

or \[-49t^2+100t-18=0\] or use \[\frac{\sqrt{b^2-4ac}}{a}\] with \[a=-49, b = 100, c = 18\]